Problems for the Adept

Lesson 1

Simple exercises

• Draw ten concenteric circles – each ring is darker than all rings outside of it

– each ring is thicker than each outer ring (use setlinewidth)

• Draw the letter A

• Draw a tombsone with a flat top (use arcto)

• Draw a spirograph star (hint: 6 “points” is easy)

• Draw a regular polygon - make a procedure that accepts number of sides as an argument

• Basic string art: Define two lines and draw lines betweent them at regular intervals along each line.

Medium/difficult exercises

• Extend the basic string art to a circle

• Arbitrary string art: Be able to do string art between two arbitrary paths (hint available)